video A population proportion is.40. A sample of size 200 will be taken and the sample proportion p will be used to estimate the population proportion. Use z-table. Round your answers to four decimal places a. what is the probability that the sample proportion will be within ±.03 of the population proportion? (Round z value in intermediate calculations to 2 decimal places.) b. what is the probability that the sample proportion will be within ±.05 of the population proportion? (Round z value in intermediate calculations to 2 decimal places.)
Answer #1

The margin of error of a single proportion = C2

a) 0.03 = /04(1-04) z,/2、

Z_{alpha /2} = 0.866

probability = P(-0.866 < Z_{alpha /2} <0.866) = P(0.866) - P(-0.866) = 0.6135

b) 0.05 = /04(1-04) z,/2、

Z_{alpha /2} = 1.44

probability = P(-1.44 < Z_{alpha /2} <1.44) = P(1.44) - P(-1.44) = 0.8501